Squids are the fastest marine invertebrates, using a powerful set of muscles to take in and then eject water in a form of jet propulsion that can propel them to speeds of over 11.5 m/s. What speed (in m/s) would a stationary 2.00 kg squid achieve by ejecting 0.115 kg of water (not included in the squid's mass) at 3.50 m/s? Neglect other forces, including the drag force on the squid.



___m/s


1
Expert's answer
2021-10-02T16:38:42-0400

Solution.

"m_s=2.00kg;"

"m_w=0.115kg;"

"v_w=3.50m\/s;"

"(m_s+m_w)v=m_sv_s-m_wv_w;"

"m_sv_s=m_wv_w\implies v_s=\dfrac{m_wv_w}{m_s};"

"v_s=\dfrac{0.115\sdot3.50}{2.00}=0.2m/s;"

Answer: "v_s=0.2m\/s."